OpenStudy (anonymous):

PLEASE HELP ME The volume for a rectangular prism is given by the formula V = l · w · h, where l is the length of the prism, w is the width of the prism, and h is the height of the prism. If the volume of a rectangular prism with a length and width of 2 inches is 24 cubic inches, what is the height of the rectangular prism? 6 inches 12 inches 20 inches 16 inches

OpenStudy (anonymous):

thanks @sammixboo by the way thank you so much for helping me with the old question

OpenStudy (nirmalnema):

24=2*2* height 24=4*height height = 24/4 = 6inch

OpenStudy (anonymous):

thanks

OpenStudy (nirmalnema):

welcome.. :)

sammixboo (sammixboo):

The formula to find the height of a rectangular prism with the given length, width, and volume is h = \(\large \rm \frac{v}{wl}\) If the length ( l ) and width ( w ) is 2 inches we can plug in the numbers in. Also same with the volume ( v ) h = \(\large \rm \frac{24}{2~*2}\)

sammixboo (sammixboo):

SO first solve 2 * 2 :P Tell me what 2 * 2 is?